OpenStudy (thesmartestderpalive):

Solve the word problem. The formula s = gives the side length s of a regular hexagon with perimeter P. What is the side length of a regular hexagon with perimeter 15 inches? A. 2.5 in. B. 9 in. C. 21 in. D. 90 in.

zepdrix (zepdrix):

Hey, the formula didn't show up s=?

OpenStudy (thesmartestderpalive):

s=p/6

OpenStudy (thesmartestderpalive):

@zepdrix

zepdrix (zepdrix):

So s=p/6 and they gave you p=15 just plug it in and simplify. s=15/6 s=?

OpenStudy (thesmartestderpalive):

I did, I just divided that and got 2.5, but thanks! @zepdrix

zepdrix (zepdrix):

oh cool :)